the circle passes through the points (0, 0), (O, 5), (3, 3). Find the equation of the circle.
Answer by lwsshak3(11628)

The circle passes through the points (0, 0), (O, 5), (3, 3). Find the equation of the circle.
**
Standard form of equation for circle: (x-h)^2+(y-k)^2=r^2, (h,k) being the center of the circle with radius=r.
..
3 equations from coordinates of given 3 points (0, 0), (0, 5), (3,3)
1) (0-h)^2+(0-k)^2=r^2
2) (0-h)^2+(5-k)^2=r^2
3) (3-h)^2+(3-k)^2=r^2
..
solving for k
1) (0-h)^2+(0-k)^2=r^2
2) (0-h)^2+(5-k)^2=r^2
subtract eq2 fm eq1
k^2-(5-k)^2=0
k^2-(25-10k+k^2)=0
k^2-25+10k-k^2=0
25-10k=0
10k=25
k=2.5 or 5/2
..
solving for h
2) (0-h)^2+(5-k)^2=r^2
3) (3-h)^2+(3-k)^2=r^2
subtract eq3 fm eq2
h^2+(5-2.5)^2-[(3-h)^2+(3-2.5)^2]=0
h^2+(2.5)^2-[9-6h+h^2+.5^2]=0
h^2+6.25-9+6h-h^2-.25=0
6.25-9.25+6h=0
-3=-6h
h=.5 or 1/2
..
solving for r
1) (0-h)^2+(0-k)^2=r^2
(1/2)^2+(5/2)^2=r^2
1/4+25/4=r^2=26/4=13/2
..
Equation of circle:
(x-1/2)^2+(y-5/2)^2=13/2
see graph below as visual check on answer
note: a more precise algebriac check can be made by plugging in coordinates of given points into equation
..
y=±(13/2-(x-.5)^2)^.5+2.5